OpenStudy (anonymous):

3. The measures of two of the sides of an equilateral triangle are 3x + 15 in. and 7x – 5 in. What is the measure of the third side in inches?

OpenStudy (anonymous):

my answer was 30

OpenStudy (anonymous):

its an equilateral triangle right? so all the sides are the same 3x+15=7x-5

OpenStudy (anonymous):

was i wrong ?

OpenStudy (anonymous):

first off did you know how i got this ?

OpenStudy (anonymous):

what u wrote right now?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Yup yup

OpenStudy (anonymous):

can u just give me the answer i did all of the work i just want to know so i can go back and do my work over again

OpenStudy (anonymous):

okies can you solve that

OpenStudy (anonymous):

ok 5

OpenStudy (anonymous):

dang it i knew i had it right but i thought i was missing numbers so i kept doing it different ways

OpenStudy (anonymous):

but thanks for your help

OpenStudy (anonymous):

just remember the properties of the triangles

OpenStudy (anonymous):

probably its good to write it down as notes or something :P

OpenStudy (anonymous):

on the other question with the triangle was it 12 or 14 because those were the only options that i thought were right.

OpenStudy (anonymous):

for what number?

OpenStudy (anonymous):

#2

OpenStudy (anonymous):

ah you just work it out the same way its just algebra it was just a careless mistake ( i make them alot more than you do)

OpenStudy (anonymous):

but you know how to do the math right?

OpenStudy (anonymous):

yeah i got 12?

OpenStudy (anonymous):

okies just as long as you know how it works there will be mistakes (no such thing as perfection) just try to be more careful

OpenStudy (anonymous):

ok but what did you get and how i plugged the number into X

OpenStudy (anonymous):

okies (-28) +5x= 3x+28 +(-28 ) additive inverse (-5x) +(-28) + 5x = 3x + (- 5x) additive inverse -28 =3x+(-5x) simplify -1/2 [-28 = -2x ] multiply by reciprocal 14 = x

OpenStudy (anonymous):

3 side of triangle = 180 in 3x + 15= 7x-5 x= 5, if x= 5, that makes two of them equal@ 30 in 180-30-30= 120 in for the last side?

OpenStudy (anonymous):

its a given isosceles triangle no calculation is needed

OpenStudy (anonymous):

just calculation of the angles

OpenStudy (anonymous):

ok so i had the right answer like i told you but did to different procedures >.<`

OpenStudy (anonymous):

two*

OpenStudy (anonymous):

but did you understand why i did what i did?

OpenStudy (anonymous):

whats the answer...

OpenStudy (anonymous):

answer is 14

OpenStudy (anonymous):

hey i turned it in and on #3 i was right it was 30 but i put in 5 so i got that one wrong :( all the other ones were right tho :)

OpenStudy (anonymous):

....

OpenStudy (anonymous):

thanks for your help now i have to do another assignment hopefully its a little less difficult

OpenStudy (anonymous):

i dont really know how it could be 30

OpenStudy (anonymous):

umm but sorry though :(

OpenStudy (anonymous):

can you ask your teacher about that question?

OpenStudy (anonymous):

i object to that answer properties of equilateral triangles given that all angles congruent all sides congruent 3x+15=7x-15 unless my computation was incorrect ( a careless mistake) the answer should be correct as 5

OpenStudy (icalibear):

lol thanks guys

OpenStudy (anonymous):

The correct answer was 30. haha I chose 5 -.- wish I would've gone with 30 >:( 1.c 2.A 3.D 4.A 5.B :)

OpenStudy (anonymous):

@sageyy was right! Thank you so much <3